OpenStudy (anonymous):

What is the length of segment AB? 5 units 6 units 8 units 10 units

rebeccaxhawaii (rebeccaxhawaii):

wheres the graph

OpenStudy (mathmale):

The coordinates of A and B are given. Which formula would best suit finding the distance between A and B?

OpenStudy (anonymous):

i got 5 , but not sure

OpenStudy (mathmale):

Please, share your work. How else would others know how you got 5?

OpenStudy (anonymous):

4^2 + 3^2 = 16 + 9 = 25 Square root of 25 is 5

OpenStudy (anonymous):

Help

OpenStudy (anonymous):

hello

OpenStudy (anonymous):

\[\sqrt{16+9}=5 \]

OpenStudy (anonymous):

thats what i got but it said it was wrong

OpenStudy (anonymous):

anyone else?

OpenStudy (anonymous):

please

OpenStudy (anonymous):

let me look try 10 units

OpenStudy (anonymous):

you think so?

OpenStudy (anonymous):

how is it 10

OpenStudy (anonymous):

I guessed because 10 is the only 1 that looks right